# Magnetar field dynamics shaped by chiral anomalies and helicity

### v1

@@ I Simplified discussion of the CME in the induction equation
$\bm{B}_{t}=-\bm{r}\times\bm{\nabla}\Psi,\quad\quad\quad\bm{A}_{t}=-\bm{r}\times\bm{\nabla}\Phi,\quad\quad\quad\bm{B}_{p}=\bm{\nabla}\times\bm{A}_{t}=-\bm{r}\Delta\Phi+\bm{\nabla}\frac{\partial}{\partial r}\left(r\Phi\right),$ (17)
@@
II Initial magnetic field structure and optimal CME mode
@@
The radial direction in the crust is resolved with 200 grid points, capturing structures down to a few meters. Angular directions are discretized using a cubed-sphere grid [25], with $N_{\xi}=N_{\eta}=47$ points per patch across six patches, yielding $N_{\theta}=94$ and $N_{\phi}=188$ grid points, and resolving multipoles up to $\ell_{\mathrm{max}}\sim 70$, corresponding to angular scales of a few hundred meters. With this setup, we vary the radial wavenumber $k_{r}$ while keeping the initial average magnetic field and temperature fixed. For each configuration, we compute the predicted average ($\mu_{5}^{\mathrm{avg}}$) and maximum ($\mu_{5}^{\mathrm{max}}$) values of the initial chiral chemical potential $\mu_{5}(\bm{x})$
@@ III Detailed analysis of magnetic field evolution
For completeness, Figure 5 shows the generalized helicity (Eq. (6)) in the left panel and total energy conservation (Eq. (14)) in the right panel, both for Run F. The left panel depicts the time evolution of the total helicity $Q_{5}+\frac{\alpha}{\pi\,\hbar\,c}\,\chi_{m}$ (dashed blue), the losses due to the spin-flip term at each time step $-\Gamma_{5}\,\Delta t$ (dash-dotted orange), and the resulting axial charge $Q_{5}$ (dotted teal), which directly affects the magnetic field evolution. All quantities are dimensionless. The figure illustrates how the initial magnetic helicity generates a tiny (by 20 orders of magnitude) chiral asymmetry, $Q_{5}$, which remains nearly constant over time due to spin-flip processes that counterbalance further growth. In the absence of spin-flip processes, $Q_{5}$ would grow until it is of the order of the magnetic helicity.
@@
Moreover, the figure confirms that our simulations respect the generalized helicity conservation law (Eq. (6)): the change in total helicity is closely balanced by the spin-flip term, as predicted by theory. Minor discrepancies at early times (within the first few years) arise from numerical adjustments of the initial transient stage, and small surface helicity fluxes ($\propto\bm{E}\times\bm{A}$) caused by slight violations of the imposed boundary conditions. These surface contributions decay rapidly as the system self-adjusts over a few numerical iterations. Although reducing the timestep improves precision, it significantly increases computational cost. Nonetheless, helicity is conserved to a satisfactory degree for the purposes of this analysis.
@@
Figure 5: Generalized helicity and energy conservation. Left panel (Eq. (6)): Time evolution of the total helicity $Q_{5}+\frac{\alpha}{\pi\,\hbar\,c}\,\chi_{m}$ (dashed blue), the losses due to the spin-flip term at each time step $-\Gamma_{5}\,\Delta t$ (dash-dotted orange), and the resulting axial charge $Q_{5}$ (dotted teal). All quantities are in non-dimensional form. Right panel (Eq. (14)): Time evolution of the electromagnetic energy $E_{\mathrm{em}}$ (solid blue), the Joule dissipation $Q_{\mathrm{tot}}$ (dashed red), the Poynting flux $S_{\mathrm{tot}}$ (dashed green), the chiral energy $E_{5}$ (dotted orange-red), the total flip term $\Gamma_{5\mathrm{tot}}$ (dash-dotted purple), and the total energy (solid black).
[…]

### v2

@@ I Simplified discussion of the CME in the induction equation
$\bm{B}_{t}=-\bm{r}\times\bm{\nabla}\Psi,\quad\quad\quad\bm{A}_{t}=-\bm{r}\times\bm{\nabla}\Phi,$
$\bm{B}_{p}=\bm{\nabla}\times\bm{A}_{t}=-\bm{r}\Delta\Phi+\bm{\nabla}\frac{\partial}{\partial r}\left(r\Phi\right),$ (18)
@@
II Initial magnetic field structure and optimal CME mode
@@
The radial direction in the crust is resolved with 200 grid points, capturing structures down to a few meters. Angular directions are discretized using a cubed-sphere grid [27], with $N_{\xi}=N_{\eta}=47$ points per patch across six patches, yielding $N_{\theta}=94$ and $N_{\phi}=188$ grid points, and resolving multipoles up to $\ell_{\mathrm{max}}\sim 70$, corresponding to angular scales of a few hundred meters. With this setup, we vary the radial wavenumber $k_{r}$ while keeping the initial average magnetic field and temperature fixed. For each configuration, we compute the predicted average ($\mu_{5}^{\mathrm{avg}}$) and maximum ($\mu_{5}^{\mathrm{max}}$) values of the initial chiral chemical potential $\mu_{5}(\bm{x})$.
@@ III DETAILED ANALYSIS OF MAGNETIC FIELD EVOLUTION
For completeness, Figure 5 shows the generalized helicity (Eq. (6)) in the left panel and the total energy conservation (Eq. (15)) in the right panel, both for Run F. The left panel shows the time evolution of three quantities: the total helicity $Q_{5}+\frac{\alpha}{\pi\,\hbar\,c}\,\chi_{m}$ (dashed blue); the losses due to the spin-flip term at each time step $\Gamma_{5}\,\Delta t$ (dash-dotted orange), representing the number of right-handed electrons flipping to left-handed ones in the entire crust during one time step; and $Q_{5}$ itself (dotted teal), which reflects the total excess of right-handed electrons. The figure illustrates how the magnetic helicity generates an extremely small $Q_{5}$ (20 orders of magnitude smaller), which remains nearly constant. This is due to the strong damping effect of the spin-flip term. In the absence of spin-flip processes, $Q_{5}$ would grow until it is of the order of the magnetic helicity. Moreover, the figure confirms that our simulations respect the generalized helicity conservation law (Eq. (6)): the change in total helicity is closely balanced by the spin-flip term, as predicted by theory. Minor discrepancies at early times (within the first few years) arise from adjustments of the initial transient stage, and are caused by small surface helicity fluxes ($\propto\bm{E}\times\bm{A}$) due to numerical limitations imposing boundary conditions. These surface contributions decay rapidly as the system self-adjusts over a few iterations. Although reducing the timestep improves precision, it significantly increases computational cost. Nonetheless, helicity is conserved to a satisfactory degree for the purposes of this analysis.
@@
Figure 5: Generalized helicity and energy conservation. Left panel (Eq. (6)): Time evolution of the total helicity $Q_{5}+\frac{\alpha}{\pi\,\hbar\,c}\,\chi_{m}$ (dashed blue), the losses due to the spin-flip term at each time step $\Gamma_{5}\,\Delta t$ (dash-dotted orange), and the resulting axial charge $Q_{5}$ (dotted teal). All quantities are in non-dimensional form. Right panel (Eq. (15)): Time evolution of the electromagnetic energy $E_{\mathrm{em}}$ (solid blue), the Joule dissipation $\int Q_{\mathrm{tot}}\,dt$ (dashed red), the Poynting flux $\int S_{\mathrm{tot}}\,dt$ (dashed green), the total flip term $\int\Gamma_{5\mathrm{tot}}\,dt$ (dash-dotted purple), the chiral energy $E_{5}$ (dotted orange-red), and the total energy $E_{\mathrm{tot}}$ (solid black).
To further ensure the physical robustness of our results and rule out numerical artifacts, we analyze the total energy balance (Eq. (15)). Before proceeding, we briefly examine the evolution equations for the electromagnetic energy and the energy associated with the chiral imbalance:
$\frac{\partial E_{em}}{\partial t}$ $=$ $-Q_{\rm tot}-\int dV\frac{\alpha\mu_{5}}{\pi\hbar}\bm{E}\cdot\bm{B}-S_{\rm tot},$
$\frac{\partial E_{5}}{\partial t}$ $=$ $-\Gamma_{{5}_{\rm tot}}+\int dV\frac{\alpha\mu_{5}}{\pi\hbar}\bm{E}\cdot\bm{B}~.$ (29)
In the quasi-equilibrium regime considered here, the timescale of the spin-flip term is much shorter than that of chiral density evolution, so $\partial n_{5}/\partial t\rightarrow 0$ and hence $\partial E_{5}/\partial t\rightarrow 0$. It follows that
$\Gamma_{{5}_{\rm tot}}=\int\frac{\alpha\mu_{5}}{\pi\hbar}\bm{E}\cdot\bm{B}dV~.$ (30)
The electromagnetic energy evolution then simplifies to:
$\frac{\partial E_{em}}{\partial t}=-Q_{\rm tot}-S_{\rm tot}-\Gamma_{{5}_{\rm tot}}.$ (31)
We define the total energy, $E\mathrm{tot}$, which should remain constant over time, as:
$E_{\mathrm{tot}}=E_{\mathrm{em}}-\int Q_{\mathrm{tot}}\,dt-\int S_{\mathrm{tot}}\,dt-\int{\Gamma_{5}}_{\mathrm{tot}}\,dt.$ (32)
The right panel of Figure 5 shows the time evolution of all relevant energy components, including electromagnetic energy ($E_{\mathrm{em}}$), cumulative Joule dissipation ($\int Q_{\mathrm{tot}}\,dt$), net Poynting flux ($\int S_{\mathrm{tot}}\,dt$), the total spin-flip dissipation ($\int{\Gamma_{5}}_{\mathrm{tot}}\,dt$). Denoting a global (or total) wavenumber by $k_{\mathrm{t}ot}^{2}\equiv{\left(\bm{\nabla}\times\bm{B}\right)^{2}/B^{2}}$, $Q_{\mathrm{tot}}$ and ${\Gamma_{5}}_{\mathrm{tot}}$ can be expressed as
$Q_{\mathrm{tot}}=\frac{1}{4\pi}\int dV\eta(\bm{\nabla}\times\bm{B}-k_{5}\bm{B})^{2}\propto\left(k_{\mathrm{t}ot}^{2}+k_{5}^{2}-2k_{5}k_{\bm{B}}\right)B^{2},$ (33)
${\Gamma_{5}}_{\mathrm{tot}}=\frac{1}{4\pi}\int dV\eta(\bm{\nabla}\times\bm{B}-k_{5}\bm{B})\cdot k_{5}\bm{B}\propto\left(k_{5}k_{\bm{B}}-k_{5}^{2}\right)B^{2}.$ (34)
Notably, both expressions are structurally similar and are expected to be of comparable magnitude, differing mainly due to the distinction between the global wavenumber $k_{\mathrm{t}ot}$ and the parallel wavenumber $k_{\bm{B}}$. The $k_{5}$ corrections in Equation (33) have a significant dynamical impact: they enhance dissipation when $k_{5}k_{B}$ is negative, reduce it when positive, and can completely offset magnetic dissipation when $k_{B}=k_{5}$. Overall, the total energy $E_{\mathrm{tot}}$ remains conserved to within less than $1\%$ throughout the simulation, with gradual magnetic energy loss dominated by Ohmic dissipation.